OpenStudy (august899):

I need helping finishing this question? Given that tanθ=-4.580 where 3π/2<θ<2π,find the values of sinθ and cosθ.

OpenStudy (august899):

I already started on the problem and got sin^2 θ+cos^2 θ=1 (-4.580cosθ)^2 +cos^2 θ=1 -20.976 cos^2 θ+cos^2 θ=1 -19.976cos^2 θ=1

OpenStudy (anonymous):

lets use a right triangle approach

OpenStudy (anonymous):